OpenStudy (ciarabelen):

The volume of a cylindrical can is 32pi cubic inches. If the height of the can is 2 inches, what is the radius, in inches?

OpenStudy (ciarabelen):



OpenStudy (anonymous):

@CiaraBelen volume of the cylinder = pi * sqr(r) * h where r = radius and h = height of the can volume is given in problem use the value of volume in formula and solve for r

OpenStudy (anonymous):

remember area of the cylinder=\[\pi*r^*h\]

OpenStudy (anonymous):

\[\pi*r^2*h\]

OpenStudy (anonymous):

sorry for the 1st post now substitute the values

OpenStudy (anonymous):

\[\pi*r^2*2=32\pi\]